Scale discoloration of lizards has a complete dominance relationship where the green scales are dominant over blue scales. There are 1,000 individuals with the genotype GG, 515 individuals with the genotype Gg, and 60 individuals with the genotype gg.

Find the frequency of the dominant and recessive alleles and the frequency of individuals with dominant, heterozygous and recessive genotype.

Please show your work, thank you!
Answer #1

Hardy-Weinberg equilibrium:
p+q = 1
p2+2pq+q2 = 1

Total population = 1000+515+60 = 1575

Frequency of GG, p2 = 1000/1575 = 0.635
Frequency of G, p = 0.796

Frequency of gg, q2 = 60/1575 = 0.038
Frequency of g, q = 0.194

Frequency of Gg, 2pq = 515/1575 = 0.326